# The large scale geometry of inverse semigroups and their maximal group images

### v1

@@ 4. Distortion and (weak) $F$-Inverse properties
Let $S$ and $T$ be $F$-inverse monoids, and denote their free product by $M$. We will define a premorphism $\varphi\colon M/\sigma\to M$ such that $\sigma\circ\varphi=\mathrm{id}_{M/\sigma}$, that is, a map satisfying $\varphi(1)=1,\varphi(g^{-1})=\varphi(g)^{-1}$, $\varphi(g)\varphi(h)\leq\varphi(gh)$ and $\sigma(\varphi(g))=g$ for all $g,h\in M/\sigma$. It will then follow that $M$ is $F$-inverse: indeed, given such a $\varphi$, it follows from [22, Theorem VII.6.11] that
@@
The elements of $M/\sigma=S/\sigma\ast T/\sigma$ can be uniquely represented by products $g_{1}g_{2}\cdots g_{n}$ where $g_{i}$ belong alternately to $S/\sigma\setminus\{1\}$ or $T/\sigma\setminus\{1\}$. For any $g\in S/\sigma$ or $g\in T/\sigma$, define $g^{\mathfrak{m}}$ to be the element of $S$, respectively $T$, which is greatest in the $\sigma$-class $g$. Notice that since $M$ contains isomorphic copies of both $S$ and $T$ (see [22, Section I.13.1]), $g^{\mathfrak{m}}$ can always be interpreted in $M$. We define the map $\varphi\colon S/\sigma\ast T/\sigma\to M$ by mapping $g_{1}g_{2}\cdots g_{n}$ into $g_{1}^{\mathfrak{m}}g_{2}^{\mathfrak{m}}\cdots g_{n}^{\mathfrak{m}}$, where empty products as usual are interpreted as $1$.
[…]
@@
Denote the subset of generators occurring in the relations by $Z$, and put $Y=X\setminus Z$. Then $S$ splits as a free product of $\operatorname{FIM}(Y)$ and $T=\operatorname{Inv}\langle Z\mid w_{i}=1\ (i\in I)\rangle$. Free inverse monoids are well-known to be $F$-inverse, and $T$ is $F$-inverse by Proposition 4.10, hence the statement follows by Lemma 4.13.
@@
[…]
@@
Let $u,w\in(X\cup X^{-1})^{\ast}$ label $1\to g$ paths in $\Gamma(G)$. Then $[w]_{S}\geq[u]_{S}$ if and only if for each vertex $h$ traversed by $w$, there exists some vertex $k$ traversed by $u$ with $k\in h\overline{S\Gamma}$.
@@
[…]
@@
[…]
@@ 5. Computable distortion
Let $S$ be a finitely generated inverse semigroup with maximal group image $G$, and let $l_{S}$ and $l_{G}$ be the standard word metrics corresponding to some finite generating set of $S$ and the corresponding generating set of $G$. We say that a finitely generated inverse semigroup $S$ has recursively bounded group distortion if there exists a recursive function $\phi:\mathbb{N}\to\mathbb{N}$ such that $l_{S}(s)\leq\phi(l_{G}(s\sigma))$ for all $s\in S$.
@@
Let $H$ be a subgroup of a group $G$. Take $H^{\prime}$ to be an isomorphic copy of $H$ disjoint from $G$, and $f\colon H^{\prime}\to G$ an embedding with image $H$. Then we may consider the Clifford inverse semigroup with maximal subgroups $H^{\prime}$ and $G$ and structure morphism $f$, that is, an inverse monoid $S$ on the set $H^{\prime}\sqcup G$ with multiplication defined by
$gh=\begin{cases}\textrm{the $H^{\prime}$-product }gh&\textrm{ if }g,h\in H^{\prime};\\ \textrm{the $G$-product }gh&\textrm{ if }g,h\in G;\\ \textrm{the $G$-product }f(g)\ h&\textrm{ if }g\in H^{\prime},h\in G;\textrm{ and}\\ \textrm{the $G$-product }g\ f(h)&\textrm{ if }g\in G,h\in H^{\prime}.\end{cases}$
Then $S$ is an inverse monoid with identity element the identity of $H^{\prime}$, the identity of $G$ being the only other idempotent and the inverse of each element being its inverse in $H^{\prime}$ or $G$ as appropriate. The maximal group image is clearly $G$, with the natural morphism to it being defined by the map $f$ on $H^{\prime}$ and the identity function on $G$.
Now suppose $G$ and $H$ are both finitely generated. Choose a finite generating set $X$ for $H$ and extend it to a finite generating set $Y$ for $G$. Let $X^{\prime}=\{x^{\prime}:x\in X\}$ denote a set in bijection with $X$, which we consider a generating set for $H^{\prime}$ in the obvious way. Then $S$ is clearly generated by the set $X^{\prime}\sqcup Y$. We let $d_{S}$ denote the corresponding word metric. Observe that since $[x]_{G}=[f(x^{\prime})]_{G}$ for any $x\in X$, the word metric on $G$ is the same with respect to $Y$ as the one inherited from $S$, which we unambiguously denote by $d_{G}$. Consider furthermore the word metric $d_{H}$ on $H$ corresponding to the generating set $X$; identifying $H$ and $H^{\prime}$, this coincides with the metric $d_{S}$ on $H^{\prime}$ (but usually not with the metric $d_{S}$ on $H\subseteq G$). We now have that $\sigma\colon S\to G$ restricts to an isometry on $G$, and therefore
@@
It is straightforward to see using Stephen’s algorithm that $S\Gamma(1)$ is isomorphic to the Cayley graph of $H^{\prime}$ with respect to the generating set $X^{\prime}$, and that $S\Gamma(e)$ is isomorphic to the Cayley graph of $G$ with respect to the generating set $Y\cup\{e\}$ where $[e]_{G}=1$. Furthermore it follows from (R1) that $[ww^{-1}]_{M}=1$ for every word $w$ in the alphabet $X^{\prime}\cup(X^{\prime})^{-1}$, and similarly it follows from (R3) and (R4) that $[ww^{-1}]_{M}=[e]_{M}$ for every word $w$ in $X^{\prime}\cup(X^{\prime})^{-1}\cup Y\cup Y^{-1}$ which involves at least one letter from $Y\cup Y^{-1}$. This means that $1$ and $[e]_{M}$ are the only idempotents of $M$, which is therefore a union of its two maximal subgroups $V(S\Gamma(1))\cong H^{\prime}$ and $V(S\Gamma(e))\cong G$. Observe furthermore that identifying $M$ with $H^{\prime}\sqcup G$ and defining $f(h):=eh$ for $h\in H^{\prime}$ results by (R3) in the exact same multiplication rules for $M$ as the ones defining $S$, yielding $S\cong M$.
It is well known [7, Proposition 2.1] that if $G$ is finitely presented with decidable word problem, then the distortion of the embedding of $H$ in $G$ is bounded by a recursive function if and only if $H$ has decidable membership in $G$. Thus, by taking $G$ to be a finitely presented group having a finitely presented subgroup with undecidable membership problem (an example of which can be found in [1, Theorem 1.2]), the above construction yields an inverse monoid with group distortion which is bounded but not recursively bounded.
@@
A finitely generated inverse semigroup $S=\langle X\rangle$ is called effectively $F$-inverse if it is $F$-inverse, and there is an algorithm which on input a word $w\in(X\cup X^{-1})^{\ast}$, outputs a representative of the maximal element $[w]_{S}^{\mathfrak{m}}$.
@@
Since $S$ is effectively $F$-inverse, there is a computable function, $f$ say, which maps each word $w$ to an upper bound on $l_{S}([w]^{\mathfrak{m}})$; indeed one may compute a word representing $[w]^{\mathfrak{m}}$ and then set $f(w)$ to be the number of letters in this word, which by sub-additivity of the length function exceeds the length of the element represented.
@@
Clearly the function $\phi$ is computable, since $f$ is computable and we can enumerate the finitely many words $w$ of length less than $n$ [^4] We remark that this step may not work for length functions other than those arising from word metrics. And for every $s\in S$ we may choose a word $w$ representing $s\sigma\in G$ of length $l_{G}(s\sigma)$, whereupon
@@
As in the proof of Lemma 4.9, for each $g\in G$ if we choose $w_{g}$ to label a path from $1$ to $g$ in the (necessarily connected) graph $S\Gamma\cup gS\Gamma$, then $[w_{g}]_{S}$ is the greatest element of $\sigma^{-1}(g)$. So it suffices to show that we can find such a path. If $G$ has a decidable word problem, then in particular the ball of each radius $r$ in the Cayley graph of $G$ is effectively constructible, and for any $g$, approximating $S\Gamma\cup gS\Gamma$ by Stephen’s procedure in larger and larger portions of the group Cayley graph will eventually find some choice of $w_{g}$. The fact that $S$ has recursively bounded group distortion then follows from Proposition 5.4. ∎
@@
Let $w\in(X\cup X^{-1})^{\ast}$, and write $w$ as $w_{1}\ldots w_{n}$ where each factor $w_{i}$ belongs either to $(Y\cup Y^{-1})^{+}$ or $(Z\cup Z^{-1})^{+}$, and represents a non-identity element in $\operatorname{FG}(Y)$, respectively $H$. If $w_{i}\in(Y\cup Y^{-1})^{+}$, then a representative $u_{i}$ of $[w_{i}]_{\operatorname{FIM}(Y)}^{\mathfrak{m}}$ is the freely reduced form of $w_{i}$, which is effectively computable. If $w_{i}\in(Z\cup Z^{-1})^{+}$, then a representative $u_{i}$ of $[w_{i}]_{T}^{\mathfrak{m}}$ is computable as $T$ is effectively $F$-inverse. This yields a computable representative $u=u_{1}\ldots u_{n}$ of $[w]^{\mathfrak{m}}_{S}$, and so $S$ is effectively $F$-inverse. By Proposition 5.4, $S$ also has recursively bounded group distortion. ∎
@@
For the following result we are grateful to John Meakin who (in private communication) established a special case of it and in doing so provided the idea of the proof for the general case we state here.
@@
The inverse monoid $S$ as defined above is $F$-inverse, and has the property that every Schützenberger graph isometrically embeds into $\Gamma$ via $\sigma$ (so in particular, has computable group distortion).
@@
where $w_{i}\in(X\cup X^{-1})^{+}$, $k_{i}\in\mathbb{Z}$ and $k_{1},\ldots,k_{l}\neq 0$. Observe that since $w$ is assumed to label a geodesic path, $[w_{i}]_{G}\neq 1$ for any $i$.
Since $\Delta$ is connected, there exists a $u\to v$ path in $\Delta$, with label $w^{\prime}\in(X\cup X^{-1}\cup\{t\}\cup\{t^{-1}\})^{*}$. Then $[w^{\prime}]_{G\ast\operatorname{FG}(t)}=[w]_{G\ast\operatorname{FG}(t)}$. We can without loss of generality assume that the $w^{\prime}$ labels a simple path, i.e. $w^{\prime}$ contains no subword with value $1$ in $G\ast\operatorname{FG}(t)$. Put
@@
[…]
@@
If an $E$-unitary, special inverse monoid is $F$-inverse due to satisfying the conditions of Proposition 4.10 or 4.14 and its maximal group image has a solvable word problem, then it is also effectively $F$-inverse. We ask whether this is true for all (special) $F$-inverse monoids.
@@
If $S=\operatorname{Inv}\langle X\mid u_{1}=v_{1},\ldots,u_{n}=v_{n}\rangle$ is $F$-inverse, and the word problem in $S/\sigma$ is decidable, is there an algorithm that on input $w\in(X\cup X^{-1})^{\ast}$ outputs a representative for $[w]_{S}^{\mathfrak{m}}$? What about if $S$ is special?
The examples in Section 5 do satisfy this property: it follows from the proof of Proposition 5.8 that for any $g\in G\ast\operatorname{FG}(t)$, the greatest element of $S$ in the $\sigma$-class $g$ can be represented by any element of the form $w=t^{k_{1}}w_{1}t^{k_{2}}\ldots t^{k_{l}}w_{l}t^{k_{l+1}}$, where $[w]_{G\ast\operatorname{FG}(t)}=g$, $w_{i}\in(X\cup X^{-1})^{+}$, $k_{i}\in\mathbb{Z}$, $k_{1},\ldots,k_{l}\neq 0$, and $[w_{i}]_{G}\neq 1$ for any $i$, since such a word is always readable from $1$ in a Schützenberger graph containing $1$ and $g$. If the word problem in $G$ is solvable, than any word in $(X\cup X^{-1}\cup\{t\}\cup\{t^{-1}\})^{*}$ representing $g$ can be effectively reduced to one of the form of $w$.

### v2

@@ 4. Distortion and (weak) $F$-Inverse properties
Let $S$ and $T$ be $F$-inverse monoids, and denote their free product by $M$. We will define a premorphism $\varphi\colon M/\sigma\to M$ such that $\sigma\circ\varphi=\mathrm{id}_{M/\sigma}$, that is, a map satisfying $\varphi(1)=1,\varphi(g^{-1})=\varphi(g)^{-1}$, $\varphi(g)\varphi(h)\leq\varphi(gh)$ and $\sigma(\varphi(g))=g$ for all $g,h\in M/\sigma$. It will then follow that $M$ is $F$-inverse: indeed, given such a $\varphi$, it follows from [21, Theorem VII.6.11] that
@@
The elements of $M/\sigma=S/\sigma\ast T/\sigma$ can be uniquely represented by products $g_{1}g_{2}\cdots g_{n}$ where $g_{i}$ belong alternately to $S/\sigma\setminus\{1\}$ or $T/\sigma\setminus\{1\}$. For any $g\in S/\sigma$ or $g\in T/\sigma$, define $g^{\mathfrak{m}}$ to be the element of $S$, respectively $T$, which is greatest in the $\sigma$-class $g$. Notice that since $M$ contains isomorphic copies of both $S$ and $T$ (see [21, Section I.13.1]), $g^{\mathfrak{m}}$ can always be interpreted in $M$. We define the map $\varphi\colon S/\sigma\ast T/\sigma\to M$ by mapping $g_{1}g_{2}\cdots g_{n}$ into $g_{1}^{\mathfrak{m}}g_{2}^{\mathfrak{m}}\cdots g_{n}^{\mathfrak{m}}$, where empty products as usual are interpreted as $1$.
Observe that $\varphi(g)^{-1}=\varphi(g^{-1})$ is immediate from $(g^{\mathfrak{m}})^{-1}=(g^{-1})^{\mathfrak{m}}$, so the only condition that requires proof is $\varphi(g)\varphi(h)\leq\varphi(gh)$. If $g=g_{1}\cdots g_{n}$ and $h=h_{1}\cdots h_{m}$, then the unique representative of $gh$ is obtained from $g_{1}\cdots g_{n}h_{1}\cdots h_{m}$ by iteratively replacing consecutive letters from the same group by their products, and erasing any instances of $1$. Notice that for any $g,h\in S/\sigma\cup T/\sigma$ belonging to the same group, we have $\varphi(g)\varphi(h)=g^{\mathfrak{m}}h^{\mathfrak{m}}\leq(gh)^{\mathfrak{m}}=\varphi(gh)$, since $g^{\mathfrak{m}}h^{\mathfrak{m}}\mathrel{\sigma}(gh)^{\mathfrak{m}}$. It follows by a formal induction that $\varphi$ is a premorphism. ∎
@@
Denote the subset of generators occurring in the relations by $Z$, and put $Y=X\setminus Z$. Then $S$ splits as a free product of $\operatorname{FIM}(Y)$ and $T=\operatorname{Inv}\langle Z\mid w_{i}=1\ (i\in I)\rangle$. Free inverse monoids are well-known to be $F$-inverse, and $T$ is $F$-inverse by Proposition 4.11; hence the statement follows by Lemma 4.14.
@@
We now move on to give a necessary and sufficient condition for an $E$-unitary special inverse monoid $S$ to be $F$-inverse. First, assume that $u,w\in(X\cup X^{-1})^{\ast}$ represent elements of $S$ in the $\sigma$-class $g$. Then $w]_{S}\geq[u]_{S}$ if and only if the Schützenberger graph $S\Gamma(u)$ contains a $uu^{-1}]_{S}\xrightarrow{w}[u]_{S}$ path. Consider the embedded copy of $S\Gamma(u)$ in $\Gamma(G)$. If the set of vertices traversed by the path $1\xrightarrow{u}g$ is $1=k_{1},k_{2},\ldots,k_{n}=g$, then the embedded copy of $S\Gamma(u)$ is exactly the subgraph $\langle 1\xrightarrow{u}g\rangle\cup\bigcup_{i=1}^{n}k_{i}S\Gamma$, where $\langle 1\xrightarrow{u}g\rangle$ denotes the subgraph induced by $1\xrightarrow{u}g$. In particular, since $S\Gamma(u)$ is an induced subgraph in $\Gamma(G)$ by [25, Lemma 3.5], it contains a $uu^{-1}]_{S}\xrightarrow{w}[u]_{S}$ path if and only if each vertex of the $1\xrightarrow{w}g$ path in $\Gamma(G)$ is contained in some $k_{i}S\Gamma$.
@@
Let $u,w\in(X\cup X^{-1})^{\ast}$ label $1\to g$ paths in $\Gamma(G)$. Then $w]_{S}\geq[u]_{S}$ if and only if for each vertex $h$ traversed by $w$, there exists some vertex $k$ traversed by $u$ with $k\in h\overline{S\Gamma}$.
@@
Let $S$ be an $E$-unitary, special inverse monoid. A word $w\in(X\cup X^{-1})^{\ast}$ represents a maximum element in the $\sigma$-class $g$ if and only if for each vertex $h$ traversed by $1\xrightarrow{w}g$, the graph $h\overline{S\Gamma}$ either contains $1$ or $g$, or it is a cut set for $\Gamma(G)$ with $1$ and $g$ falling in different components.
In particular, $S$ is $F$-inverse if and only if for every $g\in G$, there exists a $1\xrightarrow{w}g$ path where $w$ satisfies the above property.
@@
A word $w\in(X\cup X^{-1})^{\ast}$ represents a maximum element in the $\sigma$-class $g$ if and only if the condition of Lemma 4.16 holds for all possible paths $1\xrightarrow{u}g$. In particular for any vertex $h$ traversed by $w$, the graph $h\overline{S\Gamma}$ should intersect any $1\to g$ path, which is by definition equivalent to the statement of the proposition. ∎
@@ 5. Computable distortion
Let $S$ be a finitely generated inverse semigroup with maximal group image $G$, and let $l_{S}$ and $l_{G}$ be the standard word metrics corresponding to some finite generating set of $S$ and the corresponding generating set of $G$. We say that a finitely generated inverse semigroup $S$ has recursively bounded group distortion if there exists a recursive function $\phi:\mathbb{N}\to\mathbb{N}$ such that for every $s\in S$ we have $l_{S}(s)\leq\phi(l_{G}(s\sigma))$ (or equivalently, $d_{S}(s,t)\leq\phi(d_{G}(s\sigma,t\sigma))$ for the respective distances for all $s,t\in S$, see Proposition 4.4).
@@
Let $H$ be a subgroup of a group $G$. Take $H^{\prime}$ to be an isomorphic copy of $H$ disjoint from $G$, and $f\colon H^{\prime}\to G$ an embedding with image $H$. Then we may consider an inverse monoid $S$ on the set $H^{\prime}\sqcup G$ with multiplication defined by
[…]
Then $S$ is an inverse monoid with identity element the identity of $H^{\prime}$, the identity of $G$ being the only other idempotent and the inverse of each element being its inverse in $H^{\prime}$ or $G$ as appropriate. The maximal group image is clearly $G$, with the natural morphism to it being defined by the map $f$ on $H^{\prime}$ and the identity function on $G$. (We remark that $S$ is an example of a Clifford semigroup; see for example [13, Section 4.2] for further discussion of these.)
Now suppose $G$ and $H$ are both finitely generated. Choose a finite generating set $X$ for $H$ and extend it to a finite generating set $Y$ for $G$. Let $X^{\prime}=\{x^{\prime}:x\in X\}$ denote a set in bijection with $X$, which we consider a generating set for $H^{\prime}$ in the obvious way. Then $S$ is clearly generated by the set $X^{\prime}\sqcup Y$. We let $d_{S}$ denote the corresponding word metric. Observe that since $x]_{G}=[f(x^{\prime})]_{G}$ for any $x\in X$, the word metric on $G$ is the same with respect to $Y$ as the one inherited from $S$, which we unambiguously denote by $d_{G}$. Consider furthermore the word metric $d_{H}$ on $H$ corresponding to the generating set $X$; identifying $H$ and $H^{\prime}$, this coincides with the metric $d_{S}$ on $H^{\prime}$ (but usually not with the metric $d_{S}$ on $H\subseteq G$). We now have that $\sigma\colon S\to G$ restricts to an isometry on $G$, and therefore
@@
It is straightforward to see using Stephen’s algorithm that $S\Gamma(1)$ is isomorphic to the Cayley graph of $H^{\prime}$ with respect to the generating set $X^{\prime}$, and that $S\Gamma(e)$ is isomorphic to the Cayley graph of $G$ with respect to the generating set $Y\cup\{e\}$ where $e]_{G}=1$. Furthermore it follows from (R1) that $ww^{-1}]_{M}=1$ for every word $w$ in the alphabet $X^{\prime}\cup(X^{\prime})^{-1}$, and similarly it follows from (R3) and (R4) that $ww^{-1}]_{M}=[e]_{M}$ for every word $w$ in $X^{\prime}\cup(X^{\prime})^{-1}\cup Y\cup Y^{-1}$ which involves at least one letter from $Y\cup Y^{-1}$. This means that $1$ and $e]_{M}$ are the only idempotents of $M$, which is therefore a union of its two maximal subgroups $V(S\Gamma(1))\cong H^{\prime}$ and $V(S\Gamma(e))\cong G$. Observe furthermore that identifying $M$ with $H^{\prime}\sqcup G$ and defining $f(h):=eh$ for $h\in H^{\prime}$ results by (R3) in the exact same multiplication rules for $M$ as the ones defining $S$, yielding $S\cong M$.
It is well known [7, Proposition 2.1] that if $G$ is finitely presented with decidable word problem, then the distortion of the embedding of $H$ in $G$ is bounded by a recursive function if and only if $H$ has decidable membership in $G$. Thus, by taking $G$ to be a finitely presented group having a finitely presented subgroup $H$ with undecidable membership problem (an example of which can be found in [1, Theorem 1.2]), the above construction yields an inverse monoid with group distortion which is bounded but not recursively bounded.
@@
A finitely generated inverse semigroup $S=\langle X\rangle$ is called effectively $F$-inverse if it is $F$-inverse, and there is an algorithm which on input a word $w\in(X\cup X^{-1})^{\ast}$, outputs a representative of the maximal element $w]_{S}^{\mathfrak{m}}$.
@@
Since $S$ is effectively $F$-inverse, there is a computable function, $f$ say, which maps each word $w$ to an upper bound on $l_{S}([w]^{\mathfrak{m}})$; indeed one may compute a word representing $w]^{\mathfrak{m}}$ and then set $f(w)$ to be the number of letters in this word, which by sub-additivity of the length function exceeds the length of the element represented.
@@
Clearly the function $\phi$ is computable, since $f$ is computable and we can enumerate the finitely many words $w$ of length less than nn. [^4] We remark that this step may not work for length functions other than those arising from word metrics. And for every $s\in S$ there exists a word $w$ representing $s\sigma\in G$ of length $l_{G}(s\sigma)$, whereupon
@@
As in the proof of Lemma 4.10, for each $g\in G$ if we choose $w_{g}$ to label a path from $1$ to $g$ in the (necessarily connected) graph $S\Gamma\cup gS\Gamma$, then $w_{g}]_{S}$ is the greatest element of $\sigma^{-1}(g)$. So it suffices to show that we can find such a path. If $G$ has a decidable word problem, then in particular the ball of each radius $r$ in the Cayley graph of $G$ is effectively constructible, and for any $g$, approximating $S\Gamma\cup gS\Gamma$ by Stephen’s procedure in larger and larger portions of the group Cayley graph will eventually find some choice of $w_{g}$. The fact that $S$ has recursively bounded group distortion then follows from Proposition 5.4. ∎
@@
Let $w\in(X\cup X^{-1})^{\ast}$, and write $w$ as $w_{1}\ldots w_{n}$ where each factor $w_{i}$ belongs either to $(Y\cup Y^{-1})^{+}$ or $(Z\cup Z^{-1})^{+}$, and represents a non-identity element in $\operatorname{FG}(Y)$, respectively $H$. If $w_{i}\in(Y\cup Y^{-1})^{+}$, then a representative $u_{i}$ of $w_{i}]_{\operatorname{FIM}(Y)}^{\mathfrak{m}}$ is the freely reduced form of $w_{i}$, which is effectively computable. If $w_{i}\in(Z\cup Z^{-1})^{+}$, then a representative $u_{i}$ of $w_{i}]_{T}^{\mathfrak{m}}$ is computable as $T$ is effectively $F$-inverse. This yields a computable representative $u=u_{1}\ldots u_{n}$ of $w]^{\mathfrak{m}}_{S}$, and so $S$ is effectively $F$-inverse. By Proposition 5.4, $S$ also has recursively bounded group distortion. ∎
@@
For the following result we are grateful to John Meakin who (in a private communication) established a special case of it and in doing so provided the idea of the proof for the general case we state here.
@@
The inverse monoid $S$ as defined above is $F$-inverse, and has the property that every Schützenberger graph isometrically embeds into $\Gamma$ via $\sigma$ (so in particular, has recursively bounded group distortion).
@@
where $w_{i}\in(X\cup X^{-1})^{+}$, $k_{i}\in\mathbb{Z}$ and $k_{1},\ldots,k_{l}\neq 0$. Observe that since $w$ is assumed to label a geodesic path, $w_{i}]_{G}\neq 1$ for any $i$.
Since $\Delta$ is connected, there exists a $u\to v$ path in $\Delta$, with label $w^{\prime}\in(X\cup X^{-1}\cup\{t\}\cup\{t^{-1}\})^{*}$. Then $w^{\prime}]_{G\ast\operatorname{FG}(t)}=[w]_{G\ast\operatorname{FG}(t)}$. We can without loss of generality assume that the $w^{\prime}$ labels a simple path, i.e. $w^{\prime}$ contains no subword with value $1$ in $G\ast\operatorname{FG}(t)$. Put
@@
where $w^{\prime}_{i}\in(X\cup X^{-1})^{+}$, $k^{\prime}_{i}\in\mathbb{Z}$ and $k^{\prime}_{1},\ldots,k^{\prime}_{l}\neq 0$. Then since $w$ and $w^{\prime}$ have the same value in $G\ast\operatorname{FG}(t)$, and $w^{\prime}_{i}]_{G}\neq 1$ for any $i$, we must have $l=j$, $k_{i}=k^{\prime}_{i}$ and $w_{i}]_{G}=[w^{\prime}_{i}]_{G}$ for all $i$. Let $\alpha_{0},\ldots,\alpha_{l}$ and $\beta_{1},\ldots,\beta_{l+1}$ be the vertices in $\Gamma$ traversed by the path $u\xrightarrow{w}v$ at the following points:
@@
If an $E$-unitary, special inverse monoid is $F$-inverse due to satisfying the conditions of Proposition 4.11 or 4.15 and its maximal group image has a solvable word problem, then it is also effectively $F$-inverse by Proposition 5.6. We ask whether this is true for all (special) $F$-inverse monoids.
@@
If $S=\operatorname{Inv}\langle X\mid u_{1}=v_{1},\ldots,u_{n}=v_{n}\rangle$ is $F$-inverse, and the word problem in $S/\sigma$ is decidable, is there an algorithm that on input $w\in(X\cup X^{-1})^{\ast}$ outputs a representative for $w]_{S}^{\mathfrak{m}}$? What about if $S$ is special?
The examples in Section 5 do satisfy this property: it follows from the proof of Proposition 5.8 that for any $g\in G\ast\operatorname{FG}(t)$, the greatest element of $S$ in the $\sigma$-class $g$ can be represented by any element of the form $w=t^{k_{1}}w_{1}t^{k_{2}}\ldots t^{k_{l}}w_{l}t^{k_{l+1}}$, where $w]_{G\ast\operatorname{FG}(t)}=g$, $w_{i}\in(X\cup X^{-1})^{+}$, $k_{i}\in\mathbb{Z}$, $k_{1},\ldots,k_{l}\neq 0$, and $w_{i}]_{G}\neq 1$ for any $i$, since such a word is always readable from $1$ in a Schützenberger graph containing $1$ and $g$. If the word problem in $G$ is solvable, than any word in $(X\cup X^{-1}\cup\{t\}\cup\{t^{-1}\})^{*}$ representing $g$ can be effectively reduced to one of the form of $w$.